ellaisyoung:

A set of numbers is shown below: {0, 0.4, 1, 2, 3} Which of the following shows all the numbers from the set that make the inequality 4x + 1 ≥ 5 true? (1 point) {2, 3} {1, 2, 3} {0, 0.4, 1} {0.4, 1}

sophia333:

So simplify your equation

sophia333:

4x+ 1 = 5x

ellaisyoung:

ok....

ellaisyoung:

and then... what do i do?

sophia333:

Lemme figure this out rq-

ellaisyoung:

ok

sophia333:

So turn 4x into 4 multiply by 0

sophia333:

which is 0-

ellaisyoung:

so the answer is 0?

sophia333:

so now you have 0 + 1 (i think im going in the right direction so @jhonyy9 cottect me if im wrong)

sophia333:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ellaisyoung so the answer is 0? \(\color{#0cbb34}{\text{End of Quote}}\) no

ellaisyoung:

so what is it?

sophia333:

well i know the answer- i just need to figre an explanation out which is hard-

sophia333:

Take it away i think i messed up here ;-;

jhonyy9:

4x +1 >=5 like a first step subtract from both sides 1 so what will get ?

ellaisyoung:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 4x +1 >=5 like a first step subtract from both sides 1 so what will get ? \(\color{#0cbb34}{\text{End of Quote}}\) ummmm 2 and 3?

jhonyy9:

4x +1 -1 >= 5-1 STEP BY STEP PLEASE

ellaisyoung:

4x +4

jhonyy9:

no 4x >=4

ellaisyoung:

oh

jhonyy9:

divide both sides by 4

jhonyy9:

\[\frac{ 4x }{ 4 }\ge \frac{ 4 }{ 4 }\]

jhonyy9:

simplifie 4 by 4 on both sides s what will get ?

ellaisyoung:

ummm 4x/4

jhonyy9:

\[\frac{ 4x }{ 4 }\ge \frac{ 4 }{ 4 }\]

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 simplifie 4 by 4 on both sides so what will get ? \(\color{#0cbb34}{\text{End of Quote}}\)

ellaisyoung:

1 and 1

jhonyy9:

will get \[x \ge1\]

jhonyy9:

on left side remain x not 1

jhonyy9:

bc you simplfie 4 by 4

jhonyy9:

on right side remain just 1

jhonyy9:

@Az opinion pls - ty

ellaisyoung:

so whats the answer?

ellaisyoung:

What is the answer>